# Intent3D: 3D Object Detection in RGB-D Scans Based on Human Intention

### v2

@@ 3.3 Dataset Statistics
Figure 3: Dataset statistics: (a) The distribution of the text lengths of all intentions; (b) Word cloud of verbs used in all intention texts; (c) Word cloud of nouns used in all intention texts; (d) Number of different verbs used for each fine-grained class; (e) Number of different nouns used for each fine-grained class.
We generate a total of 44,990 intention texts for 209 fine-grained object classes from 1,042 ScanNet scenes. The dataset includes 63,451 instances and 7,524 distinct scene-object pairs. On average, each scene contains 61 instances and 43 texts, with each scene-object pair receiving 6 intention texts.
Given that verbs and nouns play pivotal roles in expressing intentions, we conduct an in-depth analysis of these parts of speech. Within these texts, we identify a total of 1,568 different verbs and 2,894 different nouns. The greater variety of nouns compared to the number of object categories suggests that the language of intention is diverse and incorporates a wide-ranging vocabulary from human daily life. On average, the intention texts associated with each object category utilize 58 different types of verbs and 77 different types of nouns.
For a more granular understanding, we offer detailed statistics in Fig. 3. Specifically, (a) shows the distribution of text lengths; (b) presents a word cloud of the verbs found in all intention texts, excluding the template-derived verbs "want", "need", "aim", and "intend"; (c) shows a word cloud of the nouns, omitting the template-derived pronoun "I"; (d) and (e) respectively presents the number of distinct verbs and nouns used for each fine-grained class, which are organized in descending order and displayed every tenth data point for clarity.
The dataset is split into train, val, and test sets, including 35,850, 2,285, and 6,855 samples, respectively, each with disjoint scenes. Our train set comes from ScanNet’s train split, while the val and test sets are derived from its val split.
Figure 4: IntentNet: (Backbones) PointNet++ is used to compute point features, MLP is used to encode the boxes predicted by a 3D object detector, and RoBERTa is used to encode the text. (Encoder) Attention-based blocks are use for multimodal fusion. The box features are enhanced by fusing with text features. (Decoder) Point features with top-k confidence are selected as proposed queries and then updated by attention-based blocks. Several MLPs are used to linearly project the queries for following losses. (Losses) We learn to match the candidate boxes with target by $L_{bce}$. Queries are learned to notice verbs (by $L_{vPos}$), align with verbs (by $L_{vSem}$), and align with objects (by $L_{voSem}$) after being modulated by verbs. Cascaded adaptive factors are used for optimization. The colored boxes in the loss blocks represent different feature tokens.
@@ 4 Our Method
We introduce the structure of IntentNet in 4.1. Given that 3D-IG poses a challenge in multimodal reasoning, demanding concurrent engagement in 3D perception, intention comprehension, and joint supervision of multiple objectives, we incorporate Candidate Box Matching for 3D understanding (Section 4.2), Verb-Object Alignment for text insight (Section 4.3), and Cascaded Adaptive Learning for joint supervision (Section 4.4). We elaborate on our pipeline below.
@@ 4.1 Multimodal Feature Extraction
Our proposed IntentNet adopts backbones and an encoder-decoder structure for multimodal feature extraction 4.1, similar to [20, 52].
Backbones As shown in Fig. 4 (Backbones), we use the pre-trained PointNet++ [46] and RoBERTa [37] as point cloud and language backbones to extract point features $P\in\mathbb{R}^{p\times c}$ and text features $T\in\mathbb{R}^{l\times c}$ respectively. We also use a pretrained detector, GroupFree [38], to detect 3D candidate boxes as additional visual reference. These candidate bounding boxes are subsequently encoded as box features $B\in\mathbb{R}^{b\times c}$ by MLP module.
Encoder As shown in Fig.4 (Encoder), the point, box, and text features undergo N = 3 layers of attention-based blocks for multimodal fusion. Specifically, in the Point Attention Block, point features perform self-attention and then cross-attention on text and box features. In the Box Attention Block, box features are enhanced by performing cross-attention on text features for the following matching objective. In the Text Attention Block, text features first perform self-attention and then cross-attention on point features.
Decoder As shown in Fig. 4 (Decoder), similar to [20, 52], point features are projected to predict whether they are within target boxes. We select the top 50 point features, linearly project them as queries, and update them across M = 6 Decoder Attention Blocks, where queries are involved in self-attention, cross-attention on text features, cross-attention on box features and cross-attention on point features. Finally, the queries are fed into four MLPs to compute the alignment with verb and object text tokens, and predict the boxes. The box regression loss $L_{box}$ is the combination of L1 loss and 3D IoU loss, same as [20, 52].
@@ 4.2 Candidate Box Matching
Simple use of candidate box features serves as an additional visual reference for better scene understanding in previous 3D-VG methods [20, 52]. However, unlike 3D-VG which explicitly specifies the target object in its referential language, 3D-IG requires models to automatically infer the desired object in the scene. Therefore, to tackle this reasoning challenge, it is necessary to explicitly identify candidate boxes in close proximity to the target for improved 3D perception associated with the intention-related task. In Fig. 4 (Losses) Candidate Box Matching, we identify matched candidate boxes using binary cross entropy:
$\begin{split}Confidence=\text{Sigmoid}(\text{MLP}(B)),~{}L_{bce}=\text{BCE}(Confidence,GT)\end{split}$ (1)
where MLP and Sigmoid project box tokens into confidence scores, $Confidence\in\mathbb{R}^{b\times 1}$. $L_{bce}$ is the binary cross entropy. $GT$ is the label, which is set to 1 for boxes whose IoU with target boxes exceeds 0.25, and 0 otherwise.
@@ 4.3 Verb-Object Alignment
Unlike referential language, where a single noun primarily denotes the target object, intention language demands that the model concurrently comprehends all verb-object pairs to fully grasp the intended behavior (see Fig. 4 (lower left)), such as "want"–"something", "support"–"back", and "relieve"–"pressure". Evidence of this can be seen in the statistics 3.3, where each object is described by a variety of verbs and nouns. Therefore, as seen in Fig. 4 (Losses), to enhance intention understanding, we propose that the queries first develop the awareness of the verb tokens within the intention text by Verb Position Prediction, and then align with the semantics of the verb tokens by Query-Verb Contrastive Learning, and finally, when modulated by the verb tokens, align with the semantics of the corresponding object tokens by Modulated Query-Object Contrastive Learning. Specifically, we use the part-of-speech tagging and dependency parsing from spaCy [17] to extract the position of the verbs and their corresponding objects (nouns). For example, "010100" indicates a 6-token sentence with the verbs appearing at the second and the fourth position. To recognize the verbs within the intention text, the queries are fed into a MLP to predict the distribution of verbs, while the label is attained by applying softmax to the verbs’ position:
$\begin{split}\textit{V}_{dist}=\text{Softmax}(V_{pos}),~{}{L}_{vPos}=\text{CE}(\text{MLP}(Q),\textit{V}_{dist})\end{split}$ (2)
where $V_{pos}\in\mathbb{R}^{l}$ is verbs’ position, $Q$ is query, ${L}_{vPos}$ is the cross-entropy loss. Secondly, to comprehend the sematic of the verbs, we align the queries with the verb tokens by contrastive learning:
where $\mathcal{T}$ and $\mathcal{Q}$ are the set of all the text tokens and queries, which are linearly projected to the same dimension. $\mathcal{T}_{l}$ and $\mathcal{Q}_{l}$ are their subsets for the linked verb tokens and queries. Specifically, the matched queries of bipartite matching [5] is linked with the verb tokens. We append "not mentioned" at sentence ends, and link unmatched queries with the "not mentioned" token. $k$ and $lv$ are the number of queries and verb tokens. Thirdly, to further comprehend the verb-object combination, we compel the queries to infer the object token from the provided verb token. Specifically, both the queries and verb tokens undergo linear projection. For each verb-object combination in the sentence, the queries are modulated by element-wise multiplication with the verb token. Subsequently, we employ the same contrastive learning in Eq. 3 to align the modulated queries with the corresponding object token, denotes as ${L}_{voSem}$. In inference, we calculate the confidence score based on $Softmax(Q^{T}T_{verb})$, where $Q$ is query and $T_{verb}$ is verb tokens. We omit temperature and normalization in the similarity calculation in the above statement for ease of reading. We will open-source the code after acceptance to help everyone grasp the details.
@@ 4.4 Cascaded Adaptive Learning
Since we deal with multiple losses, optimizing all the objectives simultaneously becomes challenging. However, there is an inherent chain of logic among these objectives. Intuitively, the loss $L_{vPos}$ should be addressed before $L_{vSem}$, as the queries should first identify the verbs before they can understand their semantics. Similarly, $L_{vSem}$ should precede $L_{voSem}$, as understanding the semantics of the verbs is necessary before inferring the object token of that verb token. Finally, $L_{voSem}$ is a prerequisite for $L_{box}$, as understanding the verb-object combination is essential for intention understanding before intention-related detection. Inspired by Focal Loss [32], adaptively adjusting the loss based on difficulty can effectively optimize the learning process and make the best of training data. Therefore, we leverage the aforementioned logic chain to propose the Cascaded Adaptive Learning scheme that transforms a loss of higher priority into a factor to adaptively adjust the loss of the subsequent priority. Specifically, given the loss of A, say $Loss_{A}$, we compute the adaptive factor by $f(x)=x.sigmoid()+0.5$ (0.5 makes sure the factor is larger than 1). Then we module the $Loss_{B}$ which has the subsequent priority by $Loss_{B}=Loss_{B}*f(Loss_{A})$. As shown in Fig.4 (Losses), we cascade this process, i.e. $L_{vPos}\rightarrow L_{vSem}\rightarrow L_{voSem}\rightarrow L_{box}$.
5 Experiments
We conduct comprehensive evaluations, including quantitative comparisons on our val and test set 5.4, ablation studies 5.5, and provide qualitative results 5.6.
5.1 Baselines
We construct several baselines by adopting three main types of language-related 3D object detection methods for the 3D-IG challenge: two expert models, one foundation model, and one LLM-based model.
Expert model We adopt two existing 3D-VG expert methods, BUTD-DETR [20] and EDA [52] (current SOTA). These models use PointNet++ [46] and RoBERTa [37] as point cloud and language backbones. Additionally, they incorporate box candidates predicted from a pretrained detector [38] as additional visual reference and use transformer layers for multimodal fusion. Limited point features are then chosen for the decoding and prediction stages. We train these models from scratch using all ground truth boxes for all target instances.
Foundation model We fine-tune the SOTA foundation model, 3D-VisTA [62], on our benchmark. They use a pretrained 3D segmenter [49] to segment candidate objects and PointNet++ to extract each object’s point clouds. The text branch uses four layers of BERT [12]. A stack of transformer layers is used for multimodal fusion. The model undergoes pretraining across various 3D tasks with unsupervised losses. For detection, it is fine-tuned to identify the best match instances and uses boxes from segmentation as detection results.
LLM-based model Chat-3D v2 [18] is the leading LLM-based method in 3D-VG. They use a 3D segmenter [21] to identify object candidates and encode them with a pretrained SOTA 3D encoder [61]. They then project object features into the LLM space and fine-tune Vicuna-7B [7], incorporating attribute, relation, and object ID learning. Object ID strings are extracted from LLM responses and the corresponding boxes, derived from segmentation, serve as detection outputs. We follow their code’s instruction, formatting all instance annotations into question-answering scripts for training. In evaluation, we calculate the confidence score for each ID strings by applying Softmax to their logits of each token prediction and choosing the highest value as the tokens’ score. The final confidence score for each object ID is then the product of these scores.
5.2 Implementation Details
For baselines, we follow their configuration on ScanRefer [6], using their official implementations. However, for BUTD-DETR [20], we adopt the implementation in EDA [52] since we do not provide text span labeling, which has been further discussed in [52]. We train these two expert models from scratch. For the foundation model, 3D-VisTA [62], we adopt their pretrained checkpoint and finetune it. For the LLM-based model, Chat-3D-v2 [18], we also use their pretrained checkpoint and conduct both zero-shot evaluation and finetuning based on official settings. Specifically, for finetuning, we follow their setting to train the last two layers of Vicuna-7B [7] along with the rest linear layers. Our IntentNet is trained from scratch for 90 epochs with a 24 batch size. The learning rate is 0.001 for PointNet++ and 0.0001 for the rest of the network, which is decayed by 0.1 at the 65th epoch. The RoBERTa is frozen. The number of point tokens is 1024, and the maximum length for text tokens is set to 256. The hidden dimension used is 288. All the methods are evaluated on the val and test set of Intent3D using their best checkpoint based on the AP@0.5 from the val set.
@@ 5.2 Implementation Details
BUTD-DETR [20] GroupFree [38] 47.12 24.56 31.05 13.05
EDA [52] Group-Free [38] 43.11 18.91 14.02 5.00
3D-VisTA [62] Mask3D [49] 42.76 30.37 36.1 19.93
Chat-3D-v20 [18] PointGroup [21] 5.86 5.24 0.15 0.13
Chat-3D-v2 [18] PointGroup [21] 36.71 32.78 3.23 2.58
IntentNet (ours) GroupFree [38] 58.34 40.83 41.90 25.36
@@
BUTD-DETR [20] GroupFree [38] 47.86 25.74 31.41 13.46
EDA [52] Group-Free [38] 44.00 19.62 14.56 5.18
3D-VisTA [62] Mask3D [49] 43.88 31.44 37.29 22.00
Chat-3D-v20 [18] PointGroup [21] 5.63 4.93 0.14 0.11
Chat-3D-v2 [18] PointGroup [21] 33.46 29.32 2.67 2.1
IntentNet (ours) GroupFree [38] 58.92 42.28 44.01 27.60

### v3

@@ 3.3 Dataset Statistics
Figure 3: (a) The distribution of the text lengths of all intentions; (b) Word cloud of verbs used in all intention texts; (c) Word cloud of nouns used in all intention texts; (d) Number of different verbs used for each fine-grained class; (e) Number of different nouns used for each fine-grained class.
We have generated a total of 44,990 intention texts for 209 fine-grained classes from 1,042 ScanNet scenes. The dataset includes 63,451 instances and 7,524 distinct scene-object pairs. On average, each scene contains 61 instances and 43 texts, with each scene-object pair receiving 6 intention texts. Given that verbs and nouns play pivotal roles in expressing intentions, we conduct an in-depth analysis of these parts of speech. Specifically, we identify a total of 1,568 different verbs and 2,894 different nouns. The greater variety of nouns compared to the number of object categories suggests that the language of intention is diverse and incorporates a wide-ranging vocabulary from daily life. On average, the intention texts associated with each object category use 58 different types of verbs and 77 different types of nouns. For a more granular understanding, we offer detailed statistics in Fig. 3. Specifically, (a) shows the distribution of text lengths; (b) presents a word cloud of the verbs of all intention texts, excluding the template-derived verbs “want”, “need”, “aim”, and “intend”; (c) shows a word cloud of the nouns, omitting the template-derived pronoun “I”; (d) and (e) present the number of distinct verbs and nouns used for each fine-grained class in descending order, with every tenth data point displayed for clarity. The dataset is split into train, val, and test sets, containing 35850, 2285, and 6855 samples, respectively, each with disjoint scenes. Our train set comes from ScanNet’s train split, while the val and test sets are derived from its val split.
Figure 4: IntentNet: (Backbones) PointNet++ is used to extract point features, MLP encodes the boxes predicted by a 3D object detector, and RoBERTa encodes the text input. (Encoder) Attention-based blocks are used for multimodal fusion, enhancing box features through integration with text features. (Decoder) Point features with top-k confidence are selected as proposed queries and then updated by attention-based blocks. Several MLPs are used to linearly project the queries for subsequent loss calculations. (Losses) The model learns to match the candidate boxes with target objects using $L_{bce}$. Queries are trained to identify verbs ($L_{vPos}$), align with verbs ($L_{vSem}$), and align with objects ($L_{voSem}$). Cascaded adaptive factors, which hierarchically weigh each loss based on its dependency on previous losses, are used for optimization. The colored boxes in the loss blocks represent different feature tokens.
@@ 4 Our Method
We introduce the structure of IntentNet in Section 4.1. Given that 3D-IG poses a challenge in multimodal reasoning, demanding concurrent engagement in 3D perception, intention comprehension, and joint supervision of multiple objectives, we incorporate Candidate Box Matching for 3D understanding (Section 4.2), Verb-Object Alignment for text insight (Section 4.3), and Cascaded Adaptive Learning for joint supervision (Section 4.4). We elaborate on our pipeline below.
@@ 4.1 Multimodal Feature Extraction
Backbones As shown in Fig. 4 (Backbones), we use the pretrained PointNet++ [Qi et al., 2017] and RoBERTa [Liu et al., 2019] as point cloud and language backbones to extract point features $P\in\mathbb{R}^{p\times c}$ and text features $T\in\mathbb{R}^{l\times c}$ respectively. We also use a pretrained detector, GroupFree [Liu et al., 2021b], to detect 3D candidate boxes as additional visual reference. These candidate bounding boxes are subsequently encoded as box features $B\in\mathbb{R}^{b\times c}$ by MLP module.
Encoder As shown in Fig. 4 (Encoder), the point, box, and text features undergo N = 3 layers of attention-based blocks for fusion. In the Point Attention Block, point features perform self-attention and then cross-attention with text and box features. In the Box Attention Block, box features are enhanced by cross-attention with text features for the following matching objective. In the Text Attention Block, text features perform self-attention and then cross-attention with point features.
Decoder As shown in Fig. 4 (Decoder), similar to [Jain et al., 2022; Wu et al., 2023], point features are projected to predict whether they are within target boxes. We select the top 50 point features, linearly project them as queries, and update them across M = 6 Decoder Attention Blocks, where queries are involved in self-attention, cross-attention with text features, cross-attention with box features and cross-attention with point features. Finally, the queries are fed into four MLPs to compute the alignment with verb and object text tokens, and predict the boxes. The box regression loss $L_{box}$ is the combination of L1 loss and 3D IoU loss, same as [Jain et al., 2022; Wu et al., 2023].
@@ 4.2 Candidate Box Matching
Unlike 3D-VG which explicitly specifies target object in its referential language, 3D-IG requires models to automatically infer the desired object in the scene. Therefore, to tackle this reasoning challenge, it is necessary to explicitly identify candidate boxes in close proximity to the target for improved 3D perception associated with the intention. In Fig. 4 (Losses) Candidate Box Matching, matched candidate boxes are computed by minimizing the binary cross entropy (BCE):
$\begin{split}Confidence=\text{Sigmoid}(\text{MLP}(B)),~{}L_{bce}=\text{BCE}(Confidence,GT),\end{split}$ (1)
where MLP followed by Sigmoid projects box tokens $B$ into confidence scores. $L_{bce}$ is the BCE. $GT$ is the label, which is set to 1 for boxes exceeding 0.25 IoU threshold and 0 otherwise.
@@ 4.3 Verb-Object Alignment
Unlike referential language, where a single noun primarily denotes the target object, intention language demands that the model concurrently comprehends all verb-object pairs to fully grasp the intended behavior (see Fig. 4 (lower left)), such as “want”–“something”, “support”–“back”, and “relieve”–“pressure”. Evidence of this can be seen in the statistics 3.3, where each object is described by a variety of verbs and nouns. Therefore, as seen in Fig. 4 (Losses), we propose a three-step approach to enhance intention understanding. The queries develop an awareness of the verb tokens within the intention text through Verb Position Prediction. Specifically, we use part-of-speech tagging and dependency parsing from spaCy [Honnibal et al., 2020] to extract the position of the verbs and their corresponding objects (nouns). For example, “010100” indicates a 6-token sentence with the verbs appearing at the second and the fourth position. To recognize the verbs within the intention text, the queries are fed into MLP to predict the distribution of verbs, while the label is obtained by applying softmax to the verbs’ position:
$\begin{split}\textit{V}_{dist}=\text{Softmax}(V_{pos}),~{}{L}_{vPos}=\text{CE}(\text{MLP}(Q),\textit{V}_{dist}),\end{split}$ (2)
where $V_{pos}\in\mathbb{R}^{l}$ is verbs’ position, $Q$ is query, ${L}_{vPos}$ is Cross Entropy loss. The queries align with the semantics of the verb tokens using Query-Verb Contrastive Learning:
where $\mathcal{T}$ and $\mathcal{Q}$ are the set of all the text tokens and queries, which are linearly projected to the same dimension. $\mathcal{T}_{l}$ and $\mathcal{Q}_{l}$ are their subsets for the linked verb tokens and queries. The matched queries from bipartite matching [Carion et al., 2020] are linked with the verb tokens. We append “not mentioned” at the end of the sentence, and link unmatched queries with the “not mentioned” token. $k$ and $lv$ are the number of queries and verb tokens. To further comprehend the verb-object relationship, the queries need to infer the object token using the provided verb token via Modulated Query-Object Contrastive Learning. Specifically, both the queries and verb tokens undergo linear projection. For each verb-object combination in the sentence, the queries are modulated by element-wise multiplication with the verb token. Then, we employ the same contrastive learning in Eq. 3 to align the modulated queries with the corresponding object token, denotes as ${L}_{voSem}$. In inference, we get a confidence score based on $Softmax(Q^{T}T_{verb})$, where $Q$ is query and $T_{verb}$ is verb tokens. We omit temperature and normalization in the similarity calculation above for ease of reading.
@@ 4.4 Cascaded Adaptive Learning
Since we deal with multiple losses, optimizing all the objectives simultaneously becomes challenging. However, there is an inherent chain of logic among these objectives. Intuitively, the loss $L_{vPos}$ should be addressed before $L_{vSem}$, as the queries should first identify the verbs before they can understand their semantics. Similarly, $L_{vSem}$ should precede $L_{voSem}$, as understanding the semantics of the verbs is necessary before inferring the object token of that verb token. Finally, $L_{voSem}$ is a prerequisite for $L_{box}$, as understanding the verb-object combination is essential for intention understanding before intention-related detection. Inspired by Focal Loss [Lin et al., 2017], adaptively adjusting the loss based on difficulty can effectively optimize the learning process and make the best use of training data. Therefore, we use the aforementioned logical chain to create the Cascaded Adaptive Learning scheme, which turns a higher-priority loss into a factor that helps adjust the next lower-priority loss. Specifically, for a given loss, say $Loss_{A}$, we compute the adaptive factor using $f(x)=x.sigmoid()+0.5$ (0.5 makes sure the factor is larger than 1). Then, we modulate the next loss, $Loss_{B}$, by multiplying it with this factor: $Loss_{B}=Loss_{B}*f(Loss_{A})$. As shown in Fig. 4 (Losses), we cascade this process, i.e. $L_{vPos}\rightarrow L_{vSem}\rightarrow L_{voSem}\rightarrow L_{box}$.
@@
BUTD-DETR [Jain et al., 2022] GroupFree [Liu et al., 2021b] 47.12 24.56 31.05 13.05
EDA [Wu et al., 2023] GroupFree [Liu et al., 2021b] 43.11 18.91 14.02 5.00
3D-VisTA [Zhu et al., 2023] Mask3D [Schult et al., 2022] 42.76 30.37 36.1 19.93
Chat-3D-v20 [Huang et al., 2023] PointGroup [Jiang et al., 2020] 5.86 5.24 0.15 0.13
Chat-3D-v2 [Huang et al., 2023] PointGroup [Jiang et al., 2020] 36.71 32.78 3.23 2.58
IntentNet (Ours) GroupFree [Liu et al., 2021b] 58.34 40.83 41.90 25.36
@@
BUTD-DETR [Jain et al., 2022] GroupFree [Liu et al., 2021b] 47.86 25.74 31.41 13.46
EDA [Wu et al., 2023] GroupFree [Liu et al., 2021b] 44.00 19.62 14.56 5.18
3D-VisTA [Zhu et al., 2023] Mask3D [Schult et al., 2022] 43.88 31.44 37.29 22.00
Chat-3D-v20 [Huang et al., 2023] PointGroup [Jiang et al., 2020] 5.63 4.93 0.14 0.11
Chat-3D-v2 [Huang et al., 2023] PointGroup [Jiang et al., 2020] 33.46 29.32 2.67 2.1
IntentNet (Ours) GroupFree [Liu et al., 2021b] 58.92 42.28 44.01 27.60
5 Experiments
5.1 Baselines
We construct baselines by adopting three main types of language-related 3D object detection methods for the 3D-IG challenge: two expert models, one foundation model, and one LLM-based model.
Expert model We adopt two existing 3D-VG expert methods, BUTD-DETR [Jain et al., 2022] and EDA [Wu et al., 2023]. These models use PointNet++ [Qi et al., 2017] and RoBERTa [Liu et al., 2019] as point cloud and language backbones. Additionally, they incorporate box candidates predicted from a pretrained detector [Liu et al., 2021b] as additional visual reference and use transformer layers for multimodal fusion. Limited point features are then chosen for the decoding and prediction stages. We train these models from scratch.
Foundation model We fine-tune the SOTA foundation model, 3D-VisTA [Zhu et al., 2023], on our benchmark. It uses a pretrained 3D segmenter [Schult et al., 2022] to segment candidate objects and PointNet++ to extract each object’s point clouds. The text branch uses four layers of BERT [Devlin et al., 2018]. A stack of transformer layers is used for multimodal fusion. The model undergoes pretraining across various 3D tasks with unsupervised losses. For detection, it is fine-tuned to identify the best match instances and it uses boxes from segmentation as detection results.
LLM-based model Chat-3D v2 [Huang et al., 2023] uses a 3D segmenter [Jiang et al., 2020] to identify object candidates and encode them with a pretrained 3D encoder [Zhou et al., 2023]. It then projects object features into the LLM space and fine-tunes Vicuna-7B [Chiang et al., 2023], incorporating attribute, relation, and object ID learning. Object ID strings are extracted from LLM responses and the corresponding boxes, derived from segmentation, serve as detection outputs. We follow their instructions in the code, thereby formatting all instance annotations into question-answering scripts for training. During evaluation, we calculate the confidence score for each ID string by applying Softmax to the logits of each token prediction and selecting the highest value as the token’s score. The final confidence score for each object ID is the product of these token scores.
5.2 Implementation Details
For baselines, we follow their configuration on ScanRefer [Chen et al., 2020] by their official codes. However, for BUTD-DETR [Jain et al., 2022], we adopt the implementation from EDA [Wu et al., 2023] since Intent3D does not provide text span labeling, which has been further discussed in [Wu et al., 2023]. We train these two expert models from scratch. For 3D-VisTA [Zhu et al., 2023], we adopt their pretrained checkpoint and fine-tune it. For Chat-3D-v2 [Huang et al., 2023], we also use their pretrained checkpoint and conduct both zero-shot evaluation and fine-tuning based on official settings. Our IntentNet is trained from scratch for 90 epochs with a batch size of 24. The learning rate is 0.001 for PointNet++ and 0.0001 for the rest of the network, which decays by 0.1 at the 65th epoch. The RoBERTa is frozen. The number of point tokens is 1024, and the maximum length for text tokens is set to 256. The hidden dimension used is 288. All the methods are evaluated on the val and test set of Intent3D using their best checkpoint based on the AP@0.5 from the val set.